Tim mows neighborhood lawns for extra money. Suppose that he would be willing to mow one lawn for ​$11​, a second lawn for ​$18​, and a third lawn for ​$23. Also suppose that three neighbors are interested in having their lawns mowed. Mrs. Jones would be willing to pay ​$34 to have her lawn​ mowed, Mr. Wilson would be willing to pay ​$27​, and Ms. Smith would be willing to pay ​$23. If Tim offers to mow lawns for ​$23 ​each, what will be his producer​ surplus?  ​$ nothing. ​(Enter a numeric response using an​ integer.)
Answer #1

Producer Surplus = Price which producer actually charges - Price at which he is willing to supply the good

Tim charged price of 23 for each lawn

Producer Surplus from 1st lawn = 23 - 11 = 12

Producer Surplus from 2nd lawn = 23 - 18 = 5

Producer Surplus from 3rd lawn = 23 - 23 = 0

Total Producer Surplus = 12 +5 + 0 = 17